Please give step by step excel tutorialThe calculation of the coefficient of determination r2 depends on the number of independent variables. An adjusted value of r

Questions:

A) r^2=

B) r^2adj=

C) About BLANK % (number value) of the variation in y can be explained by the relationship between the variables
Answer #1

Sol:

Instal analysis tool pak in excel go to

data>data analysis>regression

Book1 Excel Tell me what you want to do . Share File Home Insert Page Layout Formulas Review View Adabala Naveenkumarvenkat s

select the data

Book1 Excel Tell me what you want to do . Share Home Insert Page Layout Fomulas Review View Adabala Naveenkumarvenkat Consoli

you i will get

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.831303
R Square 0.691065
Adjusted R Square 0.38213
Standard Error 2.83326
Observations 5
ANOVA
df SS MS F Significance F
Regression 2 35.91328 17.95664 2.236929 0.308935
Residual 2 16.05472 8.027361
Total 4 51.968
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 20.32679 22.70638 0.895202 0.465147 -77.3709 118.0244
x1 0.067788 0.123362 0.549499 0.637824 -0.463 0.598573
x2 0.115319 0.294779 0.391205 0.733389 -1.15301 1.383651

From output:

r^2=0.691065

Adjusted R Square 0.38213

C) About BLANK % (number value) of the variation in y can be explained by the relationship between the variables

0.691065=0.691065*100=69.11% variation in y is explained by model.